OpenStudy (anonymous):

Which equation will have the same solution as 3 – 5(2x + 1) = 3(3 + x)? A. 11x = –13 B. 11x = 13 C. 13x = –11 D. 13x = 11

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@heather040200

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

@hilbertboy96

OpenStudy (anonymous):

3 – 5(2x + 1) = 3(3 + x) 3-10x-5=9+3x -10x-2=9+3x -13x=11 13x=-11 Choice C